Tech Staff Meeting Notes November 5, 2015

The top priority is always repair of down stations.
NM problem stations
ET problem stations

Duty Table updated June 2014


News & Highlights


Put AO/TA conversions on hold for now. Jim requested the extra sim card now so that it could be used for several noise studies on deck. The six stations in the AO network are being decommissioned and the hardware returned to the ANSS Depot. CERI offered to take over O&M for two: U40A (near Mountain Home, AR) and X40A (east of Hot Springs). Jim did a recon on U40A and its doable but will take a little work. This effort will likely take about 6 months.

2015 Preventive Maintenance84% complete
NM BB/SM & SP: 67 done & 10 to go
NM SM: 9 done & 8 to go
AG BB/SM: 6 done & 0 to go
GPS: 13 done & 0 to go
Misc: 6 done & 2 to go

GSAR is back online at Gosnell. The school is now on the Arkansas State Network similar to AG.

freewave serial reset installed at 10 of 11 broadbands. GLAT remains.

We have 11 episensors cannabalized from the K2's. All 11 RT-147's in the NM broadbands are replaced. Dataless seed is updated.


Taken from:

ARPT dropping at night (needs more solar)
CHNM - ditto (new transmitters are not as efficient as the old)
FLPT - ditto

Misc Problems
HCAR -spikes
TYAR- spikes
CPAR -spikes

All ISIS3 hardware except that slated for PLAR is installed. Have about 26 remaining, mostly MKTA and some NMAD outer stations.

RDST has a very large pole right next to the vault that will corrupt any data during strong ground motion. It needs to get moved; Dave has a couple options on the Rhodes campus (the name will change when its moved). We have a Basalt almost ready to go so we and install new rather than move old.

NODE Tasks

The win7 upgrades at nodes (mudworm is done).


Move to Wilder is complete.

Recover mast from Richardson.

There are some inconsistencies in the configuratons of the aironets. Steve will review and sync configs.

The Luxora link was getting considerable interference causing degraded signal and connection loss. Jim changed frequencies and it got better. Consider this as a candidate for a different radio.




The waveserver continues to have occasional very gappy periods most likely due to missed messages. No errors or issues were round in logs. This will get a Win7 upgrade to see if that will correct the problem.

AZTN/HDBR: Shahram submitted a proposal to TDOT to replace all the sensors with Episensors (including new cabinets, etc). We asked USGS to provide new radios so we can have a continuous buffer of all channels at AZTN. We'll then allow access to NSMP so they can extract event data. Unofficial word is TDOT approved the funding; waiting on signed contract.


Need to review and update the power systems in 218.


Replacement station for GTTN and RBNC being evaluated and permissioned for Southern Highlands Science Learning Center at Purchase Knob, a part of the Great Smoky Mountains National Park. Need to do a noise study.


Taken from:

CLTN.HH* and CLTN.HN*-sporadic

Misc Problems
CPRT.HNN see image
VHTN.EHN see image (needs a "new" S13)
VHTN.EHE see image (needs a "new" S13)

John replaced the lengthy list below with a prioritized spreadsheet HERE .


Improve RCGA AC power and perform maintenance work on shed.

Replace BCRT Strong Motion Vault and instrument. Improve transient suppression/g rounding.

Replace GMG RT-147.

Move PMTN due to excessive noise and vandalism. The trees at Lone Mountian are t oo tall to get a signal out. Looking for an alternate site. Can we use a cell modem?


HPNC node is up. Need to earthworm commands (e.g. status, restart etc) from the ssh command line; currently remote desktop is required to interface with earthworm. Need to try to get 3 complete 3-c stations operating: MGNC, GFM and SMNC. Export those channels along with picks and migrate Mt Gibbs node (BRBC) helicorder to HPNC.

Upgrade all High Peak Node Stations from L4-3D's to S-13's. At least one bad component in L4-3's at CCNC, SMNC, TVNC, TRYN.

GFM needs to be upgraded to ISIS4. Need to do a noise study for a closer site, then move it in the spring along with upgrading to S-13.


Replace VHTN E-W and N-S Instruments.

WMTN pit needs to be moved due to erosion.

CCNC is not going through fgtn and instead repeats through ASTN. The 950MHz shot would work so it could be converted to digital.

We're still repeating CPRT through GTTN. The new internet host at the Corryton Senior Center for CPRT has been verbally approved by Greg Hill the center coordinator and the Knox County facility supervisor.


No problems at this time.

AG Network

HHAR needs a new router and power supply.

WHAR replace RT147 with episensor. Done.

FCAR AC power upgrade was abandoned. A tree was removed improving sunlight but still needs additional solar.


Needs some clean up.

The FDL wiki is here:


The seisnet group on umdrive is at You will need your uuid and password.